125 g of lithium metal reacts with 3.0 L of 2.5 mol/L sulfuric acid solution.
a)Calculate the mass of lithium remaining after the reaction is over.
b) Calculate the volume of hydrogen gas produced, if this reaction occurs at STP.
1
Expert's answer
2012-12-28T06:20:28-0500
Reaction:
2Li + H2SO4 = Li2SO4 + H2

First seek theamount of substance for the lithium and sulfuric acid:
n(Li) = 125/7 = 17.857 mol
n(H2SO4) = 3 x 2.5 = 7.5 mol

Next seek the amount ofsubstance lithium remaining after the reaction is over:
n(Li) = 17.857 - 15 = 2.857 remaining after the reaction is over.

Тhe amount of substance ofhydrogen will be:
n(H2) = n(H2SO4) = 15 mol
V(H2) = n(H2) x Vm(H2)
V(H2) = 15 mol x 22.4 mol/l = 336 liters